Exercise 7-9

Anna Garden recently opened her own basketweaving studio. She sells finished baskets in addition to selling the raw materials needed by customers to weave baskets of their own. Unfortunately, owing to space limitations, Anna is unable to carry all varieties of kits originally assembled and must choose between two basic packages.

The Basic Kit includes undyed, uncut reeds (with dye included) for weaving one basket. This basic package costs Anna $16 and sells for $30. The second kit, called Stage 2, includes cut reeds that have already been dyed. With this kit the customer need only soak the reeds and weave the basket. Anna produces the Stage 2 kit by using the materials included in the Basic Kit. Because she is more efficient at cutting and dying reeds than her average customer, Anna is able to produce two Stage 2 kits in one hour from one Basic Kit. (She values her time at $18 per hour.) The Stage 2 kit sells for $36.

Partially correct answer. Your answer is partially correct. Try again.

Prepare an incremental analysis for the Anna’s basketweaving studio. (Enter negative amounts using either a negative sign preceding the number e.g. -45 or parentheses e.g. (45).)

Sell
(Basic Kit)
Process Further
(Two Stage 2 Kits)
Net Income
Increase
(Decrease)
Sales per unit $

Entry field with incorrect answer

$

Entry field with correct answer

$

Entry field with incorrect answer

Costs per unit
   Direct materials $

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

   Direct labor

Entry field with correct answer

Entry field with incorrect answer

Entry field with incorrect answer

      Total $

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

Net income per unit $

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with correct answer
Answer #1
Incremental Analysis
Sell Process further Net Income increase
Sales per unit 30 72 42
Costs per unit
Direct material 16 16 0
Direct labor 0 18 -18
Total 16 34 -18
Net Income per unit 14 38 24

2 Sell А BC 1 Incremental Analysis Process further Net Income increase 3 Sales per unit 30 =36*2 =C3-B3 4 Costs per unit 5 Di